OpenStudy (horsegal244):

Ill Give Medals! And Fan! A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represents 60 miles? Part B How many miles is represented by 1/4 inch on the map?

OpenStudy (aaronandyson):

Really? x'D

OpenStudy (horsegal244):

What?

OpenStudy (aaronandyson):

You just helped a person with the same question.

OpenStudy (horsegal244):

No i did not....

OpenStudy (aaronandyson):

You had replied to it.

OpenStudy (horsegal244):

Yea my question....

OpenStudy (aaronandyson):

Nevertheless, A) 2 inches=25 miles x inches = 60 miles B) 2 inches=25 miles 1/4 inches= x miles

OpenStudy (horsegal244):

@Skielerlucas04

OpenStudy (horsegal244):

Can u just tell me how to do it please

OpenStudy (aaronandyson):

If 2 inches = 25 miles Than x inches = 60 miles Therefore, Inches Miles 2 25 x 60 Cross multiply 2*60 = 25*x Solve for x :)

OpenStudy (horsegal244):